The peroxydisulfate ion (S2O82-) reacts with the iodide ion in aqueous solution via the reaction:

S2O82- (aq) + 3I- → 2SO42- + I3- (aq)

An aqueous solution containing 0.050 M of S2O82-ion and 0.072 M of I-is prepared, and the progress of the reaction followed by measuring [I-].The data obtained is given the table below.

Time (s) 0.000 400.0 800.0 1200.0 1600.0
[I-] (M) 0.072 0.057 0.046 0.037 0.029

The concentration of S2O82- remaining at 800 s is ________ M.

1) 0.015
2) 0.041
3) 4.00 ⋅ 10-3
4) 0.046
5) 0.076
Answer #1

It has to:

[I-] 800 s = 0.072 - 3 * X = 0.046

It clears X = 0.009 M

It has to:

[S2O8-2] 800 s = 0.05 - X = 0.05 - 0.009 = 0.041 M

Option 2.

If you liked the answer, please rate it positively, you would help me a lot, thanks.